OpenStudy (princeevee):

solve for c a=1/4(b+c) A. c = a – 4b B. c = 4a – 4b C. c = 4a – b D. c = 4b – a

OpenStudy (welshfella):

multiply both sides of the formula by 4 what do you get/

OpenStudy (welshfella):

what does 4 * 1/4 give you?

OpenStudy (anonymous):

1 whole

OpenStudy (welshfella):

right so multiply left side of the formula by 4 gives 4a and left side becomes b + c 4a = b + c

OpenStudy (anonymous):

b+c=4a

OpenStudy (welshfella):

now its easy to rearrange this to get the correct answer

OpenStudy (welshfella):

just subtract b from both sides and you have it

OpenStudy (princeevee):

c?